OpenStudy (anonymous):

8th grade math please help!! Suppose you want to join a video store. Big Video offers a special discount card that costs $9.99 for one year. With the discount card, each video rental costs $2.49. A discount card from Main Street Video costs $20.49 for one year. With the Main Street Video discount card, each video rental costs $1.79. After how many video rentals is the cost the same?

OpenStudy (anonymous):

Suppose you are thinking about buying one of two cars. Car A will cost $17,655. You can expect to pay an average of $1230 per year for fuel, maintenance, and repairs. Car B will cost about $15,900. Fuel, maintenance, and repairs for it will average about $1425 per year. After how many years are the total costs for the cars the same?

OpenStudy (anonymous):

Well....write small little equations. Let the number of years be Y. What is the cost of car A? What is the cost of car B? Set them equal and evaluate for y.

OpenStudy (anonymous):

C=17,655+1230x B=15,900+1425x set these equal to each other and solve for x 17,655+1230x=15,900+1425x